# In-plane Black-hole Spin Measurements Suggest Most Gravitational-wave Mergers Form in Triples

### v1

@@ Abstract
The spin-orbit tilt angles $\theta_{1(2)}$ of merging stellar-mass black holes provide key insights into their astrophysical origin. GWTC4pop report that the spin-orbit tilt distribution of mergers in the latest Gravitational-Wave Transient Catalog 4.0 exhibits a global peak at near-perpendicular directions $\cos\theta_{1(2)}\approx 0$. Here, we recover this feature using hierarchical Bayesian inference with parametric models that are tailored to enhance the diagnostic power about astrophysical formation channels. We find that the spin distribution of the low-mass bulk of the binary black hole merger population ($m_{1}\lesssim 44.3^{+8.7}_{-4.6}\,\rm M_{\odot}$) can be well-modelled by a dominant Gaussian component that peaks at $\cos\theta_{1(2)}\approx 0$, possibly mixed with a subdominant isotropic component. Models that include a component with spins preferentially aligned with the orbit are disfavoured by current data (with Bayes factors $|\Delta\ln\mathcal{B}|\approx 1$ to $3$) and constrain its contribution to be small ($\xi\sim\mathcal{O}(1)\,\%$). If these findings are reinforced by more detections, they would challenge any major contribution from the traditional isolated-binary formation scenario yielding closely aligned spins. Instead, the dominant component with near-perpendicular spins qualitatively matches expectations from the evolution of isolated massive stellar triples in the galactic field, where the Lidov–Kozai effect naturally produces a unique overabundance of mergers with $\cos\theta_{1(2)}\approx 0$.
software: GWpopulation (Talbot2019; Talbot2025); Bilby (Bilby); GWpopulation Pipe (gwpop_pipe)
@@ 1 Introduction
A decade after the first direct detection of gravitational waves from merging binary black holes (Abbott2016), the observational sample detected by the LIGO-Virgo-KAGRA (LVK) interferometers has grown to about two hundred of these events (2015CQGra..32g4001L; 2015CQGra..32b4001A; 2013PhRvD..88d3007A; 2012CQGra..29l4007S; 2018LRR....21....3A; 10.1093/ptep/ptaa125; GWTC4Updating). Yet, a central question remains open: What are the formation mechanisms behind the observed mergers of black holes?
@@
Resolving this question has become a major challenge in gravitational-wave astronomy as most proposed binary black hole formation channels—such as those arising from isolated binary-star evolution (2016ApJ...819..108B), active galactic nuclei (Stone2017), or ultra-wide binaries perturbed by the Galaxy (2024ApJ...972L..19S)—are notoriously difficult to model in a predictive way (e.g., 2018PhRvD..98h4036G; Baibhav2024). Their outcomes depend on numerous uncertain assumptions, making their predictions highly flexible; with sufficient tuning, they can be made to reproduce a wide range of features in gravitational-wave observations.
@@
Among all proposed formation channels, the evolution of triples—which are by far the most abundant observed configuration of massive black hole progenitor stars (Moe2017; Offner2023)—stands out by making a unique, testable prediction for the spin orientation of merging black holes. In this scenario, the gravitational perturbation from a distant companion drives large-amplitude eccentricity oscillations (the “Lidov–Kozai” effect) of black holes formed in the inner binary (Zeipel1910; Lidov1962; Kozai1962; Naoz2016). Efficient gravitational-wave emission during close pericentre passages can then lead the binary to inspiral and merge (Silsbee2017; Antonini2017; Grishin2018; Liu2018; Antonini2018; Rodriguez2018; Stegmann2022; 2025A&A...699A.272V; Dorozsmai2025; Stegmann2025). Under well-defined conditions, the combined action of Lidov–Kozai oscillations, gravitational-wave emission, and relativistic spin precession drives the component spins to flip into the orbital plane (Antonini2018; Liu2018; Rodriguez2018), producing an overall excess of systems with spin–orbit misalignments near perpendicular (cf. Section 4). This configuration is so unusual that other formation channels typically require additional, ad hoc assumptions to reproduce it (e.g., StegmannAntonini; Tauris2022; Baibhav2024; Vaccaro2024). The triple scenario therefore provides a robust formation pathway with a clear, falsifiable prediction.
@@
Tentative evidence for a global peak in the binary black hole population at $\cos\theta\approx 0$—corresponding to a spin–orbit tilt angle of about $90^{\circ}$—was identified in analyses of earlier LVK catalogues (Vitale2022; 2023ApJ...946...16E; 2023PhRvD.108j3009G). Support for this feature has strengthened with the Gravitational-Wave Transient Catalog 4.0 (GWTC-4.0; GWTC2data; GWTC3data; GWTC4data; GWTC4Updating; GWTC4pop, see Figure 7 in the latter), particularly in the non-parametric (“weakly modelled”) B-spline population model (2023ApJ...946...16E), which makes minimal a priori assumptions, but can be difficult to interpret astrophysically. The LVK’s parametric (“strongly modelled”) default spin-tilt model (GWTC4pop) also recovers weak evidence for this peak. Here we aim to assess and interpret this feature using a more astrophysically motivated parametric framework, that allows direct comparison with predictions from formation channels.
2 Methods
@@
While many studies have shown that the population spin–tilt distribution is a powerful discriminator of black hole formation channels (2017CQGra..34cLT01V; 2017MNRAS.471.2801S; 2017PhRvD..96b3012T; Vitale2022), tilt measurements for individual events carry large uncertainties (2014PhRvL.112y1101V; 2020PhRvR...2d3096P; 2021PhRvD.104j3018B), limiting population-level inference (2024PhRvD.109j4036M; VitaleMould2025). Population constraints have therefore often relied on the effective spin $\chi_{\rm eff}={\chi_{1}\cos\theta_{1}+q\,\chi_{2}\cos\theta_{2}}/{1+q},$ where $\chi_{1(2)}$ and $\theta_{1(2)}$ denote the component spin magnitudes and tilt angles, respectively, and $0<q\leq 1$ is the binary mass ratio (Damour2001). While $\chi_{\rm eff}$ is more precisely measured, its inferred population distribution peaks near zero (2020ApJ...895..128M; 2021PhRvD.104h3010R; Banagiri2025b; GWTC4pop). This is intrinsically ambiguous: it can arise from small spin magnitudes or from substantial spin-orbit misalignment, and it is further compounded by the mass-weighted combination of both components. This motivates inferring the population distribution of individual spin magnitudes and tilts directly, rather than relying on $\chi_{\rm eff}$ alone. In what follows, we therefore employ a hierarchical Bayesian inference method (e.g., Mandel2019) to infer hyper-parameters describing the component spin properties of the binary black hole merger population.
@@
We use the GWPopulation code (Talbot2019; Talbot2025) and the public detector sensitivity estimates (2025PhRvD.112j2001E; GWTCsensitivity) and individual event posteriors from GWTC-4.0 (GWTC2data; GWTC3data; GWTC4data). We use the posterior samples obtained with the NRSur7dq4 waveform model (2019PhRvR...1c3015V) where available for new events in GWTC-4.0, the mixed-waveform samples otherwise, and the IMRPhenomXPHM waveform samples (2021PhRvD.103j4056P; 2025PhRvD.111j4019C) for all events that appear in previous catalogs. Our method is exemplified in a public script accompanying this work (StegmannCodeInPrep) and detailed in the following.
We introduce two distribution functions
@@
where $\mathcal{N}_{[a,b]}(x|\mu,\sigma)$ is a truncated Gaussian distribution within $a\leq x\leq b$ and $\mu$ and $\sigma$ are its mean and standard deviation, respectively. Using Equations (1) and (2) we define
$\pi=$ $\,(1-\zeta)\Big[\xi\,\pi_{t}(\cos\theta_{i}|\mu_{t},\sigma_{t})\,\pi_{\chi}(\chi_{i}|\mu_{\chi},\sigma_{\chi})$
$+\frac{1-\xi}{4}\,\pi_{\chi}(\chi_{i}|\mu^{\rm Iso}_{\chi},\sigma^{\rm Iso}_{\chi})\Big]$
$+\frac{\zeta}{4}\,\pi_{\chi}(\chi_{i}|\mu^{\rm HighIso}_{\chi},\sigma^{\rm HighIso}_{\chi}),$ (3)
@@
Table 1: Overview of parametric spin models studied in this work. The truncated normal distribution $\mathcal{N}_{[0.1,4]}$ used as a prior for $\sigma_{t}$ in most models is assuming a mean and standard deviation of $\mu=0$ and $\sigma=1/2$, respectively. The two LVK models enforce that the Gaussian and isotropic components are following the same spin magnitude distribution, i.e., $\mu_{\chi}^{\rm Iso}$ and $\sigma_{\chi}^{\rm Iso}$ in Equation (3) are replaced by $\mu_{\chi}$ and $\sigma_{\chi}$, respectively. Other model specifications are detailed in Section 2. The last three columns contain the excess fractions defined in Equations (5) and (6) for $\delta=0.1$ and resulting Bayes factors relative to the Gaussian + Isotropic + Cut model, which are discussed in Section 3.
@@
Equation (3) is used to investigate population models that reflect merger contributions from various astrophysical formation channels. An isotropic component is expected from binary black hole mergers formed through close few-body encounters in dense environments such as star clusters (e.g., Rodriguez2016). Mergers from the evolution of isolated binary and triple systems are expected to have preferentially aligned (Kalogera2000; Farr2017; 2018PhRvD..98h4036G; Belczynski2020; Bavera2020; OlejakSpins2021; Broekgaarden2022) or near-perpendicular spin-orbit orientations (Antonini2018; Liu2018), respectively, and can be reflected by the Gaussian component with appropriate $\mu_{t}$ and $\sigma_{t}$ (cf. Section 4). Meanwhile, mergers assembled in active galactic nucleus disks are less certain but generally expected to inherit a preferred axis set by the disk angular momentum (Yang2019a; Vaccaro2024).
@@
Thus, we introduce a set of parametric models whose properties are summarised in Table 1 and priors on the population parameters are defined at the end of this section. In the Gaussian and Aligned models we assume that the spin-orbit tilts can be described by a single truncated Gaussian whose location is allowed to be free within $1\leq\mu_{t}\leq 1$ or enforced at alignment $\mu_{t}=1$, respectively. In these single-component models we impose $\zeta=0$ and $\xi=1$. In the next set of models Gaussian + Isotropic, Uniform Gaussian + Isotropic, and Aligned + Isotropic we allow for a free mixing with an isotropic component by allowing $0\leq\xi\leq 1$. In these models, the Gaussian and isotropic components are allowed to follow different spin magnitude distributions each parametrised by truncated Gaussians with $(\mu_{\chi},\sigma_{\chi})$ and $(\mu_{\chi}^{\rm Iso},\sigma_{\chi}^{\rm Iso})$, respectively. If both components are thought to represent two different astrophysical subpopulations, e.g., from the evolution of isolated stellar few-body systems (binaries, triples, etc.) and from few-body encounters in dense environments (e.g., globular clusters), respectively, there is a priori no reason to assume they should follow the same spin magnitude distribution. This is different from the default modelling of GWTC4pop, which enforces the same spin magnitude distribution in both components. To test the consequences of this assumption, we include two models LVK Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic, where we replace $(\mu_{\chi}^{\rm Iso},\sigma_{\chi}^{\rm Iso})$ by $(\mu_{\chi},\sigma_{\chi})$.
@@
Finally, 2025PhRvL.134a1401A; Antonini2025 identify a transition to an isotropic and high-spin population for primary masses above a mass threshold $\tilde{m}=45^{+6}_{-4}M_{\odot}$, which they interpret as evidence for a high-mass tail produced through hierarchical black hole mergers (e.g., Rodriguez2019; GerosaFishbach). Consistently, 2022ApJ...941L..39W identify a similar transition to a higher-spin population at ${46.1}_{-5.1}^{+5.6}\,{M}_{\odot}$ based on GWTC–3 data. Here, we focus on the lower-mass bulk of the astrophysical merger population, but ensure that our analysis is not contaminated by the distinct spin properties of the high-mass population. Therefore, we introduce a Gaussian + Isotropic + Cut model, in which
$\zeta(m_{1},\tilde{m})=\frac{1}{1+\exp(-m_{1}+\tilde{m})}$ (4)
@@
is a sigmoid function that smoothly transitions from the Gaussian + Isotropic model at low primary masses $m_{1}\lesssim\tilde{m}$ to an isotropic component with separate spin magnitude distribution $(\mu_{\chi}^{\rm HighIso},\sigma_{\chi}^{\rm HighIso})$ at high masses $m_{1}\gtrsim\tilde{m}$. Another model Aligned + Isotropic + Cut additionally enforces $\mu_{t}=1$.
@@
In the Uniform Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic models we also assume a uniform prior $\mathcal{U}_{[0.1,4]}$ for $\sigma_{t}$, similar to population models of GWTC4pop. In all other models we assume $\mathcal{N}_{[0.1,4]}(\mu=0,\sigma=1/2)$, as wide spin tilt distributions (e.g., reflected in a uniform prior up to $\sigma_{t}=4$) are inconsistent with most astrophysical models, which instead tend to exhibit more pronounced peaks and are better modelled by a width of $\sigma\approx 1/2$ or less (e.g., Baibhav2024, and astrophysical models in Figure 2 below).
Figure 2: Posterior predictive distribution (PPD) of the black hole spin tilts of the low-mass population ($m_{1}\lesssim\tilde{m}=44.3^{+8.7}_{-4.6}\,\rm M_{\odot}$) in our parametric Gaussian + Isotropic + Cut model (blue). The PPD is constructed by taking the quantiles (median indicated by dashed lines, 90% interval by shaded envelope) across MCMC samples (thin lines). The left panel shows the PPD of only the low-mass Gaussian component, the right panel includes the low-mass isotropic component. The green colour shows the non-parametric B-Spline model of GWTC4pop; GWTC4pop-zenodo. The red line shows simulation outcomes of binary black hole mergers which are caused by the Lidov–Kozai effect in hierarchical triples (Antonini2018). The orange lines show simulation outcomes of binary black hole mergers from isolated binary star evolution (Olejak2024) assuming high natal kicks at black hole formation that are drawn from a Maxwellian velocity distribution with $\sigma=133\,\rm km\,s^{-1}$ (dash-dotted) or natal kicks lowered by fallback (Fryer2012) with $\sigma=265\,\rm km\,s^{-1}$ (solid). Since we cut the y-axis, the small inset (linear axes) shows that the orange models are strongly concentrated at $\cos\approx 1$. In the right panel, the binary black hole sketch depicts a merger with $\cos\theta_{1(2)}\approx 0$, where, for visualisation purposes, we pick near-opposite in-plane directions (which is generally poorly constrained from the data).
3 Results
At first, we quantify the statistical significance of the global peak at $\cos\theta\approx 0$ in the non-parametric B-Spline model of GWTC4pop. We follow a similar approach as Vitale2022 and define

### v2

# Gravitational-wave Observations Suggest Most Black Hole Mergers Form in Triples

@@ Abstract
The spin-orbit tilt angles $\theta_{1(2)}$ of merging stellar-mass black holes provide key insights into their astrophysical origin. Non-parametric population modelling of The LIGO, Virgo, and KAGRA Collaborations (2025c) shows that the spin-orbit tilt distribution of mergers in the latest Gravitational-Wave Transient Catalog 4.0 exhibits a global peak at near-perpendicular directions $\cos\theta_{1(2)}\approx 0$. Here, we recover this feature using hierarchical Bayesian inference with parametric models that are tailored to enhance the diagnostic power about astrophysical formation channels. We find that the spin distribution of the low-mass bulk of the binary black hole merger population ($m_{1}\lesssim 44.3^{+8.7}_{-4.6}\,\rm M_{\odot}$) can be well-modelled by a dominant Gaussian component that peaks at $\cos\theta_{1(2)}\approx 0$, possibly mixed with a subdominant isotropic component. Models that include a component with spins preferentially aligned with the orbit are disfavoured by current data (with Bayes factors $|\Delta\ln\mathcal{B}|\approx 1$ to $3$) and constrain its contribution to be likely small ($\xi\sim\mathcal{O}(1)\,\%$), although large contributions cannot yet ruled out with certainty. If these findings are reinforced by more detections, they would challenge any major contribution from the traditional isolated-binary formation scenario yielding closely aligned spins. Instead, the dominant component with near-perpendicular spins matches expectations from the evolution of isolated massive stellar triples in the galactic field, where the Lidov–Kozai effect naturally produces a unique overabundance of mergers with $\cos\theta_{1(2)}\approx 0$.
software: GWpopulation (Talbot et al., 2019, 2025); Bilby (Ashton et al., 2019); GWpopulation Pipe (Talbot, 2021)
@@ I Introduction
A decade after the first direct detection of gravitational waves from merging binary black holes (Abbott et al., 2016), the observational sample detected by the LIGO-Virgo-KAGRA (LVK) interferometers has grown to about two hundred of these events (Aasi et al., 2015; Acernese et al., 2015; Aso et al., 2013; Somiya, 2012; Abbott et al., 2018; Akutsu et al., 2020; The LIGO, Virgo, and KAGRA Collaborations, 2025e). Yet, a central question remains open: What are the formation mechanisms behind the observed mergers of black holes?
@@
Resolving this question has become a major challenge in gravitational-wave astronomy as most proposed binary black hole formation channels—such as those arising from isolated binary-star evolution (Belczynski et al., 2016), active galactic nuclei (Stone et al., 2017), or ultra-wide binaries perturbed by the Galaxy (Stegmann et al., 2024)—are notoriously difficult to model in a predictive way (e.g., Gerosa et al., 2018; Baibhav and Kalogera, 2024). Their outcomes depend on numerous uncertain assumptions, making their predictions highly flexible; with sufficient tuning, they can be made to reproduce a wide range of features in gravitational-wave observations.
@@
Among all proposed formation channels, the evolution of triples—which are by far the most abundant observed configuration of massive black hole progenitor stars (Moe and Di Stefano, 2017; Offner et al., 2023)—stands out by making a unique, testable prediction for the spin orientation of merging black holes. In this scenario, the gravitational perturbation from a distant companion drives large-amplitude eccentricity oscillations (the “Lidov–Kozai” effect) of black holes formed in the inner binary (Zeipel, 1909; Lidov, 1962; Kozai, 1962; Naoz, 2016). Efficient gravitational-wave emission during close pericentre passages can then lead the binary to inspiral and merge (Silsbee and Tremaine, 2017; Antonini et al., 2017; Grishin et al., 2018; Liu and Lai, 2018; Antonini et al., 2018; Rodriguez and Antonini, 2018; Mangipudi et al., 2022; Stegmann et al., 2022a; Vigna-Gómez et al., 2025; Dorozsmai et al., 2025; Stegmann and Klencki, 2025). Under well-defined conditions, the combined action of Lidov–Kozai oscillations, gravitational-wave emission, and relativistic spin precession drives the component spins to flip into the orbital plane (Antonini et al., 2018; Liu and Lai, 2018; Rodriguez and Antonini, 2018; Su et al., 2021), producing an overall excess of systems with spin–orbit misalignments near perpendicular (cf. Section IV). This configuration is so unusual that other formation channels typically require additional, ad hoc assumptions to reproduce it (e.g., Stegmann and Antonini, 2021; Tauris, 2022; Baibhav and Kalogera, 2024; Vaccaro et al., 2024). The triple scenario therefore provides a robust formation pathway with a clear, falsifiable prediction.
@@
Tentative evidence for a global peak in the binary black hole population at $\cos\theta\approx 0$—corresponding to a spin–orbit tilt angle of about $90^{\circ}$—was identified in analyses of earlier LVK catalogues through parametric modelling in Vitale et al. (2022) and was subsequently corroborated with more agnostic spin models (Edelman et al., 2023; Golomb and Talbot, 2023). Support for this feature has strengthened with the Gravitational-Wave Transient Catalog 4.0 (GWTC-4.0; The LIGO and Virgo Collaborations, 2022; The LIGO, Virgo, and KAGRA Collaborations, 2023a, 2025b, 2025e, 2025c, see Figure 7 in the latter), particularly in the non-parametric (“weakly modelled”) B-spline population model (Edelman et al., 2023), which makes minimal a priori assumptions, but can be difficult to interpret astrophysically. The LVK’s parametric (“strongly modelled”) default spin-tilt model (The LIGO, Virgo, and KAGRA Collaborations, 2025c) also recovers weak evidence for this peak. Here we aim to assess and interpret this feature using a more astrophysically motivated parametric framework, that allows direct comparison with predictions from formation channels.
II Methods
@@
While many studies have shown that the population spin–tilt distribution is a powerful discriminator of black hole formation channels (Vitale et al., 2017; Stevenson et al., 2017; Talbot and Thrane, 2017; Vitale et al., 2022), tilt measurements for individual events carry large uncertainties (Vitale et al., 2014; Pratten et al., 2020; Biscoveanu et al., 2021), limiting population-level inference (Miller et al., 2024; Vitale and Mould, 2025). Population constraints have therefore often relied on the effective spin $\chi_{\rm eff}={\chi_{1}\cos\theta_{1}+q\,\chi_{2}\cos\theta_{2}}/{(1+q)},$ where $\chi_{1(2)}$ and $\theta_{1(2)}$ denote the component spin magnitudes and tilt angles, respectively, and $0<q\leq 1$ is the binary mass ratio (Damour, 2001). While $\chi_{\rm eff}$ is more precisely measured, its inferred population distribution peaks near zero (Miller et al., 2020; Roulet et al., 2021; Banagiri et al., 2025; The LIGO, Virgo, and KAGRA Collaborations, 2025c). This is intrinsically ambiguous: it can arise from small spin magnitudes or from substantial spin-orbit misalignment, and it is further compounded by the mass-weighted combination of both components. This motivates investigating the population distribution of individual spin magnitudes and tilts directly, rather than relying on $\chi_{\rm eff}$ alone. In what follows, we therefore employ a hierarchical Bayesian inference method (e.g., Mandel et al., 2019) to infer hyper-parameters describing the component spin properties of the binary black hole merger population.
@@
We use the GWPopulation code (Talbot et al., 2019, 2025) and the public detector sensitivity estimates (Essick et al., 2025; The LIGO, Virgo, and KAGRA Collaborations, 2025a) and individual event posteriors from GWTC-4.0 (The LIGO and Virgo Collaborations, 2022; The LIGO, Virgo, and KAGRA Collaborations, 2023a, 2025b). In using GWPopulation, we limit the variance of the Monte Carlo–estimated log-likelihood to be below 1 and use the posterior samples obtained with the NRSur7dq4 waveform model (Varma et al., 2019) where available for new events in GWTC-4.0, the mixed-waveform samples otherwise, and the IMRPhenomXPHM waveform samples (Pratten et al., 2021; Colleoni et al., 2025) for all events that appear in previous catalogs. Our method is exemplified in a public script [^1] https://github.com/stegmaja/black-hole-spin-orbit-tilts accompanying this work and detailed in the following.
In order to streamline our notation, we introduce two helper functions
@@
where $\mathcal{N}_{[a,b]}(x|\mu,\sigma)$ is a truncated Gaussian distribution within $a\leq x\leq b$ and $\mu$ and $\sigma$ are its mean and standard deviation, respectively. Using Equations (1) and (2) we define the distribution function
${\pi(\chi_{i},\cos\theta_{i}|\Lambda)}=(1-\zeta)\Big[\xi\,\pi_{t}(\cos\theta_{i}|\mu_{t},\sigma_{t})\pi_{\chi}(\chi_{i}|\mu_{\chi},\sigma_{\chi})$
$+\frac{1-\xi}{4}\,\pi_{\chi}(\chi_{i}|\mu^{\rm Iso}_{\chi},\sigma^{\rm Iso}_{\chi})\Big]+\frac{\zeta}{4}\,\pi_{\chi}(\chi_{i}|\mu^{\rm HighIso}_{\chi},\sigma^{\rm HighIso}_{\chi}),$ (3)
${\{\zeta,\xi,\mu_{t},\sigma_{t},\mu_{\chi},\sigma_{\chi},\mu^{\rm Iso}_{\chi},\sigma^{\rm Iso}_{\chi},\mu^{\rm HighIso}_{\chi},\sigma^{\rm HighIso}_{\chi}\}\in\Lambda}$
@@
Table 1: Overview of parametric spin models studied in this work. The truncated normal distribution $\mathcal{N}_{[0.1,4]}$ used as a prior for $\sigma_{t}$ in most models is assuming a mean and standard deviation of $\mu=0$ and $\sigma=1/2$, respectively. The two LVK models enforce that the Gaussian and isotropic components are following the same spin magnitude distribution, i.e., $\mu_{\chi}^{\rm Iso}$ and $\sigma_{\chi}^{\rm Iso}$ in Equation (3) are replaced by $\mu_{\chi}$ and $\sigma_{\chi}$, respectively. Other model specifications are detailed in Section II. The last three columns contain the excess fractions defined in Equations (5) and (6) for $\delta=0.1$ and resulting Bayes factors relative to the Gaussian + Isotropic + Cut model, which are discussed in Section III.
@@
Equation (3) is used to investigate population models that reflect merger contributions from various astrophysical formation channels. An isotropic component is expected from binary black hole mergers formed through close few-body encounters in dense environments such as star clusters (e.g., Rodriguez et al., 2016). Mergers from the evolution of isolated binary and triple systems are expected to have preferentially aligned (Kalogera, 2000; Farr et al., 2017; Gerosa et al., 2018; Belczynski et al., 2020; Bavera et al., 2020; Olejak and Belczynski, 2021; Broekgaarden et al., 2022) or near-perpendicular spin-orbit orientations (Antonini et al., 2018; Liu and Lai, 2018), respectively, and can be reflected by the Gaussian component with appropriate $\mu_{t}$ and $\sigma_{t}$ (cf. Section IV). Meanwhile, mergers assembled in active galactic nucleus disks are less certain but generally expected to inherit a preferred axis set by the disk angular momentum (Yang et al., 2019; Vaccaro et al., 2024).
@@
Thus, we introduce a set of parametric models whose properties are summarised in Table 1 and priors on the population parameters are defined at the end of this section. In the Gaussian and Aligned models we assume that the spin-orbit tilts can be described by a single truncated Gaussian whose location is allowed to be free within $1\leq\mu_{t}\leq 1$ or enforced at alignment $\mu_{t}=1$, respectively. In these single-component models we impose $\zeta=0$ and $\xi=1$. In the next set of models Gaussian + Isotropic, Uniform Gaussian + Isotropic, and Aligned + Isotropic we allow for a free mixing with an isotropic component by allowing $0\leq\xi\leq 1$. In these models, the Gaussian and isotropic components are allowed to follow different spin magnitude distributions each parametrised by truncated Gaussians with $(\mu_{\chi},\sigma_{\chi})$ and $(\mu_{\chi}^{\rm Iso},\sigma_{\chi}^{\rm Iso})$, respectively. If both components are thought to represent two different astrophysical subpopulations, e.g., from the evolution of isolated stellar few-body systems (binaries, triples, etc.) and from few-body encounters in dense environments (e.g., globular clusters), respectively, there is a priori no reason to assume they should follow the same spin magnitude distribution. This is different from the default modelling of The LIGO, Virgo, and KAGRA Collaborations (2025c), which enforces the same spin magnitude distribution in both components. To test the consequences of this assumption, we include two models LVK Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic, where we replace $(\mu_{\chi}^{\rm Iso},\sigma_{\chi}^{\rm Iso})$ by $(\mu_{\chi},\sigma_{\chi})$.
@@
Finally, Antonini et al. (2025a, b) identify a transition to an isotropic high-spin population for primary masses above a mass threshold $\tilde{m}=45^{+6}_{-4}M_{\odot}$, which they interpret as evidence for a high-mass tail produced through hierarchical black hole mergers (e.g., Rodriguez et al., 2019; Gerosa and Fishbach, 2021). Consistently, Wang et al. (2022) identify a similar transition to a higher-spin population at ${46.1}_{-5.1}^{+5.6}\,{M}_{\odot}$ based on GWTC–3 data. Here, we focus on the lower-mass bulk of the astrophysical merger population, but ensure that our analysis is not contaminated by the distinct spin properties of the high-mass population. Therefore, we introduce a Gaussian + Isotropic + Cut model, in which
$\zeta(m_{1},\tilde{m})=\frac{1}{1+\exp[\tilde{k}(\tilde{m}-m_{1})]}$} (4)
@@
is a sigmoid function that smoothly transitions from the Gaussian + Isotropic model at low primary masses $m_{1}\lesssim\tilde{m}$ to an isotropic component with separate spin magnitude distribution $(\mu_{\chi}^{\rm HighIso},\sigma_{\chi}^{\rm HighIso})$ at high masses $m_{1}\gtrsim\tilde{m}$. Another model Aligned + Isotropic + Cut additionally enforces $\mu_{t}=1$. As a default, we fix $\tilde{k}=1$ in all our models which incorporate the mass cut. We have also tested one variant of the Gaussian + Isotropic + Cut model, where we allow $\tilde{k}$ to be a free hyper-parameter of the model between 0.1 and 2.0 (where larger values of $\tilde{k}$ lead to a sharper transition in mass). As a result, we found that $\tilde{k}=0.9^{+0.7}_{-0.6}$ is poorly constrained with a posterior distribution which is largely consistent with the prior on $\tilde{k}$, which we assumed to be uniform. As shown in Appendix B, we also found all other posterior distributions of the model parameters to be largely similar between the runs with free and a fixed $\tilde{k}=1$.
@@
In the Uniform Gaussian + Isotropic and LVK Uniform Gaussian + Isotropic models we also assume a uniform prior $\mathcal{U}_{[0.1,4]}$ for $\sigma_{t}$, similar to population models of The LIGO, Virgo, and KAGRA Collaborations (2025c). In all other models we assume $\mathcal{N}_{[0.1,4]}(\mu=0,\sigma=1/2)$, as wide spin tilt distributions (e.g., reflected in a uniform prior up to $\sigma_{t}=4$) are inconsistent with most astrophysical models, which instead tend to exhibit more pronounced peaks and are better modelled by a width of $\sigma\approx 1/2$ or less (e.g., Baibhav and Kalogera, 2024, and astrophysical models in Figure 2 below).
Figure 2: Posterior predictive distribution (PPD) of the black hole spin tilts of the low-mass population ($m_{1}\lesssim\tilde{m}=44.3^{+8.6}_{-4.6}\,\rm M_{\odot}$) in our parametric Gaussian + Isotropic + Cut model (blue). The PPD is constructed by taking the quantiles (median indicated by dashed lines, $90\,\%$ interval by shaded envelope) across MCMC samples (thin lines). The left panel shows the PPD of only the low-mass Gaussian component, the right panel includes the low-mass isotropic component. The green colour shows the non-parametric B-Spline model of The LIGO, Virgo, and KAGRA Collaborations (2025c, d). In both panels, the dotted blue lines indicate the prior bounds ($90\,\%$), i.e., they show the PPDs if all hyperparameters are drawn from their priors. The red line shows simulation outcomes of binary black hole mergers which are caused by the Lidov–Kozai effect in hierarchical triples (Antonini et al., 2018). The orange lines show simulation outcomes of binary black hole mergers from isolated binary star evolution (Olejak et al., 2024) assuming high natal kicks at black hole formation that are drawn from a Maxwellian velocity distribution with $\sigma=133\,\rm km\,s^{-1}$ (dash-dotted) or natal kicks lowered by fallback (Fryer et al., 2012) with $\sigma=265\,\rm km\,s^{-1}$ (solid). Since we cut the y-axis, the small inset (linear axes) shows that the orange models are strongly concentrated at $\cos\approx 1$. In the right panel, the binary black hole sketch depicts a merger with $\cos\theta_{1(2)}\approx 0$, where, for visualisation purposes, we pick near-opposite in-plane directions.
III Results
At first, we quantify the statistical significance of the global peak at $\cos\theta\approx 0$ in the non-parametric B-Spline model of The LIGO, Virgo, and KAGRA Collaborations (2025c). We follow a similar approach as Vitale et al. (2022) and define